OpenStudy (thcsbabygirl1224):

Which function has a phase shift pi/2 of to the right? y = 2sin( x +pi ) y = 2sin(2x +pi ) y = 2sin(x +pi ) y = 2sin(x −pi ) y = 2sin(2x −pi )

OpenStudy (thcsbabygirl1224):

@Hyuna301

OpenStudy (thcsbabygirl1224):

@gottennis121 @desmarie @truwhovian11 @misty1212

OpenStudy (aveline):

The phase shift is equal to -C/B

OpenStudy (thcsbabygirl1224):

im lost sorry what is -c/b

OpenStudy (aveline):

y = A sin(Bx + C) + D

OpenStudy (aveline):

\[\frac{ \pi }{ 2 }=\frac{ -C }{ B }\]